### v2

[…]
Grace Period is All You Need: Individual Fairness without Revenue Loss in Revenue Management
[…]
Imagine you and a friend purchase identical items at a store, yet only your friend received a discount. Would your friend’s discount make you feel unfairly treated by the store? And would you be less willing to purchase from that store again in the future? Based on a large-scale online survey that we ran on Prolific, it turns out that the answers to the above questions are positive. Motivated by these findings, in this work we propose a notion of individual fairness in online revenue management and an algorithmic module (called “Grace Period”) that can be embedded in traditional revenue management algorithms and guarantee individual fairness. Specifically, we show how to embed the Grace Period in five common revenue management algorithms including Deterministic Linear Programming with Probabilistic Assignment (DLP-PA), Resolving Deterministic Linear Programming with Probabilistic Assignment (RDLP-PA), Static Bid Price Control (S-BPC), Booking Limit (BL), and Nesting (NESTING), thus covering both stochastic and adversarial customer arrival settings. Embedding the Grace Period does not incur additional regret for any of these algorithms. This finding indicates that there is no tradeoff between a seller maximizing their revenue and guaranteeing that each customer feels fairly treated.
1 Introduction
Revenue management plays a pivotal role in various sectors including airline, retail, advertising, and hospitality, as evidenced by previous research Williamson (1992), Talluri and Van Ryzin (1998, 2004), Talluri et al. (2004). In the most standard version, sequential decision-making for revenue management encompasses the following seller-customer interaction: The seller has access to limited-capacity resources of potentially more than one types. Customers of different types arrive sequentially (the order of arrival can be either stochastic or adversarial) over a finite period of time. Upon arrival, the seller can choose either to accept (i.e., sell the item to them) or reject (i.e., not sell) the customer. Acceptance leads to revenue generation equal to the payment made by the customer, alongside the consumption of specific units of resources based on the customer’s type. On the other hand, rejection of a customer does not lead to any revenue or resource usage. The goal of the seller is to maximize the total anticipated revenue throughout the given period, while adhering to the capacity constraints of all resources.
This work emphasizes the equitable treatment of customers through the requirement of individual fairness. The concept of fairness is crucial not only from an ethical standpoint but also for its impact on customer behavior and potential to increase revenue. For example, think about the following (hypothetical, for now) scenario. Two stores, A and B, offer an identical product at a same price. However, Store A selectively offers discounts leading to perceived unfairness among customers. One may think that such perceived unfair treatment may deter customers from buying at Store A, possibly benefiting Store B. Okun (2011) explains that fair treatment to each individual is instrumental for a profit-maximizing firm to the maximization of long-run profits. To confirm if customer reactions align with our aforementioned intuition, we conducted a large-scale online survey on Prolific, detailed in Section 7. Our survey’s findings validate our priors; customers do indeed perceive various forms of unfair treatment from sellers and this in turn affects the probability of them buying again from certain stores.
We next define individual fairness in the context of revenue management.
Definition 1.1 (Individual Fairness in Revenue Management)
Let $n$ be the number of customer types, with each type $i\in[n]$ having $n_{i}$ customers arriving. An algorithm $\mathcal{A}$ is $(\alpha,\delta)$-fair for any choice of $\alpha,\delta\in(0,1)$ if for each customer of type $i$, with a probability at least $1-\delta$, for any $u,v\in[n_{i}]$,
$\mathbb{P}\left[i(v)\succ_{\mathcal{A}}i(u)\right]\leq\alpha|u-v|,$
where $i(u),i(v)$ denote the $u$-th and $v$-th customer’s of type $i$, and $i(v)\succ_{\mathcal{A}}i(u)$ denotes that algorithm $\mathcal{A}$ accepts customer $i(v)$ but not $i(u)$.
Intuitively speaking, an algorithm satisfying individual fairness ensures that the probability of two customers of the same type that arrive closely in time but receive disparate treatments is very small. Our definition of individual fairness (both the time- and the type-dependent components) is driven by our survey findings (cf. Section 7), which show that people often perceive unfairness when they do not obtain a resource that other people that are similar to them receive. As a result, we aim to upper bound the probability of customers of the same type ($i(v)$ and $i(u)$) receiving disparate treatment. As for the time component, Kahneman et al. (1986) executed a telephone survey inquiring whether customers perceived differential treatment by a hardware store during a night and the subsequent early morning, amidst a severe snowstorm, as fair. The findings indicated that 82% of respondents deemed such disparate treatment as unfair. Furthermore, our survey findings suggest that individuals perceive different treatments occurring in the near term as more unfair compared to those in the distant past. This insight suggests that the probability of two similar customers being treated differently, $\mathbb{P}\left[i(v)\succ_{\mathcal{A}}i(u)\right]$, should be moderated according to the time gap between their arrivals, specifically using the L1 norm, $\alpha|u-v|$, as a measure. Furthermore, our findings remain valid when considering measures that increase linearly or at a greater rate with respect to $|u-v|$, i.e., $\alpha|u-v|^{p}$ for any $p\geq 1$. In the existing literature, this measure is referred to as the time-variant individual fairness metric, as suggested by Gupta and Kamble (2021). A comprehensive discussion of various fairness metrics presented in prior research is available in Section 1.2.
In this paper, our primary objective is to incorporate the principle of individual fairness into the most commonly used algorithms solving the quantity-based revenue management problem, including Deterministic Linear Programming with Probabilistic Assignment (DLP-PA), Resolving Deterministic Linear Programming with Probabilistic Assignment (RDLP-PA), Static Bid Price Control (S-BPC), Booking Limit (BL), and Nesting (NESTING). We modify each algorithm one by one to meet individual fairness criteria as defined in Definition 1.1. In addition to implementing individual fairness, we also strive to maintain revenue optimization, ensuring that incorporating fairness does not lead to any loss compared to the original algorithm’s performance [^1] In Appendix 13, we show how our approach can also produce an individually-fair variant of Bid Price Control with Online Gradient Descent (BPC-OGD), albeit with a small asymptotic loss in revenue.. By adapting these well-known algorithms, we aim to ensure that any revenue management problems addressed by them also upholds the principles of individual fairness.
1.1 Main Contributions
“Grace Period” Design. In Section 3, we introduce the Grace Period which we find to be the workhorse concept for achieving individual fairness in revenue management. At a simplified level, the Grace Period works as follows: As the seller’s available resources approach depletion, an arriving customer is either accepted or rejected based on the treatment of the preceding customer. If the previous customer was accepted, the succeeding customer is likewise accepted with a probability of $(1-\alpha)$. However, if the previous customer was rejected, the incoming customer is also declined. This design (which gets utilized only when the resources are almost depleted) forms a grace period during which the probability of unequal treatment is confined. We remark that the Grace Period for customers of type $i$ can (and will in general) be different from the Grace Period for customers of type $j$ (for $i\neq j$). The tricky part is to guarantee that despite these distinct grace periods, the seller does not sell resources beyond their allowable capacity. We showcase how the Grace Period can be used via an example analysis of the First-Come-First-Serve (FCFS) algorithm.
The rationale for the Grace Period Design lies in its mitigation of unfairness. If we consider a scenario where the Grace Period Design isn’t applied and we adhere to a first-come-first-serve model, the pattern of acceptances and rejections is distinctly binary; the early arrivals are accepted, followed by a series of rejections as resources are exhausted. In such a scenario, the initial group of customers who are declined would be unfairly treated (per Definition 1.1). However, the application of the Grace Period Design spreads this perception of unfairness more equally among earlier arrivals. Hence, this design maintains individual fairness among every customer with high probability.
Grace-Period-Enhanced Algorithms for Revenue Management. In Sections 4 and 5, we integrate the Grace Period design into five well-known algorithms frequently used in revenue management (RM): DLP-PA, RDLP-PA, S-BPC, BL and NESTING. The first three DLP-PA, RDLP-PA, S-BPC are used for stochastic customer arrivals, while BL and NESTING are used for adversarial arrivals. For completeness, we outline each of these algorithms together with the explanation of why they are not individually fair in Section 2. Based on the different structure of these algorithms, we employ different methods to incorporate the grace period design into them. For example, in DLP-PA, each type $i$ customer is accepted with certain probability $p_{i}$, leading to unfairness because consecutive type $i$ customers are treated differently with a probability $p_{i}(1-p_{i})$, violating our fairness definition (1.1) for $\alpha<p_{i}(1-p_{i})$. To address this, we avoid random decisions for individual arrivals. Instead, we divide the timeline into segments, accepting a random number of customers in a first-come-first-serve manner per segment. Grace periods inside each segment ensure fairness when decisions change from acceptance to rejection within each segment; grace periods between each segment guarantee fairness when decisions change from rejection to acceptance between the segments. For clarity, we refer to the new algorithms as GP-Enhanced.
No Trade-off between Revenue Maximization and Individual Fairness in an Asymptotic Regime. Contrary to the common belief that there exists a trade-off between maximizing revenue and achieving individual fairness, we demonstrate that by augmenting standard RM algorithms with the grace period design, it is possible to attain optimal revenue (asymptotically) while simultaneously ensuring individual fairness for most of the algorithms.
For settings where the customer arrival process is stochastic, the performance of RM algorithms is measured in terms of regret with regards to the benchmark of the optimal revenue achievable (which assumes known realized demand). Table 1 contains a summary of the regret for GP-Enhanced algorithms for stochastic arrivals versus the traditional non-individually fair counterparts. For settings where the customer arrival process is adversarial, the performance of RM algorithms is measured in terms of the competitive ratio; that is, the worst-case ratio (across all possible arrival instances) between the revenue generated by the algorithm and the hindsight optimal (i.e., the maximum revenue assuming full knowledge of the realized demand). Table 2 contains a summary of the competitive ratio for GP-Enhanced algorithms for adversarial arrivals versus the traditional non-individually fair counterparts.
Table 1: Regret of Original Algorithms and Grace Period Enhanced Algorithms under Stochastic Arrival
Original Algorithm Regret of Original Algorithm Regret of Grace Period Enhanced Algorithm
DLP-PA $O(\sqrt{T})$ $O(\sqrt{T})$ (Theorem 4.1)
RDLP-PA $O(T^{1/3})$ $O(T^{1/3})$ (Theorem 4.1)
S-BPC $O(\sqrt{T})$ $O(\sqrt{T})$ (Theorem 4.2)
Table 2: Competitive Ratio of Original Algorithms and Grace Period Enhanced Algorithms under Adversarial Arrival. $C\in[0,1]$ is a setting-specific value that is specified in the analysis.
Original Algorithm Competitive Ratio of Original Algorithm Competitive Ratio of Grace Period Enhanced Algorithm
BL $C$ $C-O(\log m/m)$ (Theorem 5.1)
NESTING $C$ $C-O(\log m/m)$ (Theorem 5.2)
No Additional Computational Cost for Implementing Individual Fairness. One might conjecture that simultaneously achieving optimal revenue and being individually fair might necessitate additional computational cost for the algorithm. However, this is not the case. In all GP-Enhanced algorithms, we have preserved the algorithms’ original structure to the maximum extent possible, maintaining the order of computational complexity equivalent to that of the original algorithm.
Survey Highlighting the Significance of Individual Fairness in RM. We conducted an online survey between October and November 2023 on Prolific about people’s perception of unfair treatment in RM. Our findings suggest: (i) When customers witness behavior in a store that is considered “unfair” (as described in Definition 1.1) before they shop, their inclination to buy decreases. (ii) A customer who has bought from a store in the past and then experiences what they perceive as unfair treatment is likely to decrease their purchasing from that store in the future. (iii) Individuals perceive different treatment happening in the near term as more unfair compared to those in the distant past. With these observations, we can evaluate the importance of individual fairness to business and validate the mathematical definition of individual fairness metrics in Definition 1.1.
1.2 Other Related Work
Time Varient Individual Fairness. In the fields of Computer Science and Operations Research, individual fairness, often referenced as envy-freeness, is a critical metric in scenarios where equitable treatment of entities—such as customers, users, or resources—is desired. The time-variant aspect of individual fairness adds another layer of complexity, as it demands that customers arriving at similar times should receive similar treatment. Different papers adapt the concept of fairness to suit their specific context, which can lead to varying definitions and formulations. Here are some general axioms and mathematical formulations that capture the essence of time-variant individual fairness:
Axiom 1: Lipschitz Continuity in Time (Dwork et al. 2012, Gupta and Kamble 2021, Mukherjee et al. 2020)
This formulation states that the change in treatment should be limited by the elapsed time between arrivals, ensuring a smooth transition:
$D(f(t),f(s))\leq Ld(t,s)$
Here, $t$ and $ts$ represent the arrival time, $f(x)$ denotes the treatment function, $D$ is a distance metric measuring similarity between treatments, $d$ is a distance metric measuring the time difference, and $L$ is a Lipschitz constant that bounds the rate of change in treatment over time. This ensures that the treatment function $f$ does not vary too sharply with small changes in time.
Axiom 2: Probabilistic Fairness (Our paper), (Donahue and Kleinberg 2020, Herlihy et al. 2023)
This approach introduces randomness in treatment while maintaining fairness statistically:
$P(D\left(f(t),f(s)\right)|\leq\delta|t,s)\geq 1-\epsilon\text{ for }d(t,s)<\gamma$
where $\delta,\gamma,\epsilon$ are thresholds dictating the acceptable level of treatment difference, time difference, and probability of unequal treatment, respectively.
Axiom 3: Time-Consistent Treatment (Cohen et al. 2021, Gupta et al. 2023, Heidari and Krause 2018)
This axiom ensures that entities arriving within a short time window receive comparable treatment. Mathematically, it can be expressed as:
$\forall t,s\text{ with }d(t,s)<\epsilon,\quad D(f(t),f(s))\leq\delta$
Here, $\epsilon,\delta$ are small constants defining the permissible time window and treatment disparity, respectively.
Our fairness metric aligns with Axiom 2. Given that our time horizon is discrete, we adopt the distance measure $d(t,s)=|t-s|$. Furthermore, we have verified that substituting this with $d(t,s)=|t-s|^{p}$ for any $p\geq 1$ does not affect the validity of our results. Moreover, unlike the literature where decision variables are continuous—such as pricing or the allocation of divisible goods—our decision variable is discrete, limited to either accept or reject. Consequently, we employ $D$ as a probability measure to assess the likelihood of different treatments, which is the most appropriate measure in our context.
Mathematically speaking, Axiom 1 is stronger than Axiom 2. The intuition behind is that in Axiom 1, the treatment function must not change more rapidly than a constant rate over time. This means that the rate at which treatment changes is tightly controlled, providing a uniform smoothness in how treatments are administered over time.
Axiom 2 is stronger than Axiom 3. The intuition is that Axiom 2 generally specifies that for entities arriving within a close timeframe, the probability of receiving similar treatment must be high—usually quantified by a certain threshold. It requires a statistical guarantee that can be measured and verified. Axiom 3, while ensuring that entities arriving at similar times receive similar treatment, doesn’t specify how similar the treatments need to be or provide a statistical probability of similarity. It’s more qualitative and less stringent, making it weaker in comparison.
Individual Fairness from the Regulator’s Perspective. Individual fairness can be categorized from distinct perspectives. The first perspective considers the regulator’s perspective (Arsenis and Kleinberg 2022), suggesting that the acceptance probability for each arriving customer should be equal. The second type looks at fairness from the customer’s perspective, seeking to reduce the likelihood of individuals experiencing envy and subsequently minimize the loss associated with such envy. Our research, along with all works discussed in the previous time-variant individual fairness section, falls under this category.
These two dimensions of individual fairness differ due to their distinctive focuses. Consider an example: two customers arrive, and the first algorithm accepts both customers A and B with a probability of $0.5$. The second algorithm accepts both customers A and B with a probability of $0.1$. It appears that both algorithms uphold individual fairness from the regulator’s perspective. However, the probability that customer 1 is rejected and customer 2 is accepted under the first algorithm is $(1-0.5)\cdot 0.5=0.25$, whereas for the second algorithm, it is $(1-0.1)\cdot 0.1=0.09$. This implies that the second algorithm would potentially lead to lower feelings of envy and unfairness among customers, thus demonstrating the difference in the two metrics of individual fairness.
Modern Revenue Management Models. Due to the limitations of the traditional network revenue management model, researchers have diversified its approach for greater relevance and depth. Despite these advancements, the resolution of these contemporary models often relies on the foundational algorithmic frameworks discussed in Section 2. Our contributions aim to refine these foundational structures, thereby promoting individual fairness across all advanced revenue management models that employ decision-making algorithms outlined in Section 2. For instance, to ensure group fairness, a DLP-PA algorithm is proposed by Ma et al. (2020), while a BPC-OGD algorithm is suggested by Balseiro et al. (2021). In the context of choice-based revenue management, a choice-based DLP-PA algorithm is recommended by Liu and Van Ryzin (2008), Gallego et al. (2015). To tackle revenue management issues involving offline prediction in demand, Balseiro et al. (2022), Golrezaei et al. (2023) advocate for a BL algorithm. Revenue management with correlated arrivals is addressed by Jiang (2023) with an S-BPC method, and Jiang et al. (2020) proposes a BPC-OGD algorithm for non-stationary arrivals. In scenarios involving revenue management with reusable resources, Huo and Cheung (2022) employs a NESTING algorithm. Finally, for revenue management with horizontal uncertainty, a BPC-OGD algorithm is suggested by Balseiro et al. (2023a).
Price-based Revenue Management. Contrasting with the quantity-based model in this paper, price-based revenue management represents a distinct category. Here, pricing decisions inversely affect demand, with higher prices leading to lower demand and vice versa. Maglaras and Meissner (2006) suggests that, under certain conditions, price-based and quantity-based models align in function. Furthermore, price-based strategies also utilize the algorithmic frameworks mentioned in Section 2. Works by Gallego and Van Ryzin (1994, 1997) implement the DLP-PA algorithm for decision-making, whereas Jasin and Kumar (2012), Jasin (2014), Wang and Wang (2022) employ the RDLP-PA algorithm. In Section 8, we explore the concept of individual fairness within the price-based framework and discuss the integration of a grace period design in this context.
Group Fairness. While a considerable volume of literature exists concerning fairness in RM algorithms, much of it primarily aims to achieve group fairness (Cohen et al. 2022, Aminian et al. 2023, Balseiro et al. 2021, Cayci et al. 2020, Donahue and Kleinberg 2020, Ma et al. 2020, Manshadi et al. 2021). Group fairness in decision-making entails ensuring that every group (or type) of customer is treated equitably. However, group fairness does not guarantee individual fairness. This is because, firstly, individuals might not be aware of the groups to which they belong. Secondly, individuals within each group can be treated inequitably based on their specific contexts and features.
2 Model and Preliminaries
For ease of exposition, we present the classical network revenue management model [^2] There is a special case where the decision-maker only has one type of resource, and the special model is called as single-leg revenue management.. In the main text, we will operate under the assumption that all algorithms are applied to this classical model.
In classical network revenue management, there are $L$ types of indivisible resources, each with a capacity of $m_{j},j\in[L]$. All capacities $m_{j},\forall j\in[L]$, scale consistently, i.e., there exist constants $q_{j}$ for every $j\in[L]$ such that $m_{j}=q_{j}m$. There is a total of $T$ —where $T$ is also scaling with $m$— customers, arriving one by one. There are $n$ customer types. When a customer of type $i\in[n]$ arrives, they may request multiple units from several resource types $j$; we denote this as $A_{ij}$. Matrix $\mathbf{A}\in\mathbb{R}_{+}^{n\times L}$ is called the demand matrix. Without loss of generality, we assume that $A_{ij}\leq\bar{a}$. When a type $i$ customer arrives in the system, the decision-maker has two choices: either accept the customer, allocate $A_{ij}$ units of resource $j$, and earn $r_{i}$ in revenue, where $r_{i}$ are known constants; or decline the customer’s request. The primary objective is to maximize cumulative revenue.
We study two variants of the above model based on the customers’ arrival process. For the first variant (Section 4), we assume that customers arrive sequentially at random; i.e., each customer of type $i$ has an associated arrival rate $\lambda_{i}$. In every round, at most one customer will arrive, with the probability of a type $i$ customer arriving being $\frac{\lambda_{i}}{\sum_{i^{\prime}=0}^{n}\lambda_{i^{\prime}}}$ [^3] Typically, we normalize all arrival rates so that $\sum_{i^{\prime}=0}^{n}\lambda_{i^{\prime}}=1$.. Here, $\lambda_{0}$ represents the rate for no arrivals. In Sections 4.1, and 4.2, we study the case where the arrival rates $\lambda_{i}$ are known; Appendix 13.1 studies the case where the rates $\lambda_{i}$ are unknown. Under stochastic arrivals, regret is commonly used for measuring the performance of an online algorithm. For the second variant (discussed in Section 5), we study the case where an adversary determines both the quantity and sequence of each type of customer, commonly referred to as the worst-case/adversarial scenario. Under adversarial arrivals, the competitive ratio is the metric to study the performance of an online algorithm.
2.1 Regret and Competitive Ratio Analysis
We first focus on the stochastic arrival setting. Let $\Lambda_{i}(T)$ be the random variable corresponding to the total number of type $i$ customers arriving between $[0,T]$. The hindsight optimum denotes the optimal revenue achievable if the exact number of arrivals for each customer type is known in advance. For the classic network RM, the hindsight optimum is defined as
$\text{Rev}^{\star}=\max_{\mathbf{x}}\Big{\{}\sum_{i\in[n]}r_{i}x_{i}\quad\text{ s.t. }\sum_{i\in[n]}\mathbf{A}_{i}x_{i}\leq\mathbf{m}\text{, }0\leq x_{i}\leq\Lambda_{i}(T),\forall i\in[n]\Big{\}},$ (1)
where in Eq. (1), $\mathbf{x}$ is the decision variable vector representing the quantity of each type of customer to accept, $\mathbf{A}_{i}$ is the $i^{\text{th}}$ column of preference matrix $\mathbf{A}$, m is the vector containing the capacity of each resource $m_{j}$, and $\Lambda_{i}(T)$ is a known value representing the exact number of type $i$ arrivals.
Given that the hindsight optimum is privy to the precise count of each customer type’s arrivals, it is typically utilized as a benchmark to evaluate the online algorithms. The following definition introduces the concept of regret.
Definition 2.1
Given $\text{Rev}(\pi)$ as the expected revenue generated by the online algorithm $\pi$, the regret of algorithm $\pi$ is defined as $\text{Regret}(\pi):=\mathbb{E}[\text{Rev}^{\star}]-\text{Rev}(\pi)$.
Note that since the hindsight optimum is informed of the precise number of each type’s arrivals, the resulting revenue is an upper bound for the revenue of any online algorithm. Hence, $\text{Regret}\geq 0$, and the smaller the regret, the better.
Next, we consider the adversarial arrival setting. Let $\mathcal{I}$ represent the set of all potential arrival instances. For any $I\in\mathcal{I}$, let $\text{OPT}(I)$ denote the maximum revenue achievable for arrival instance $I$. We denote $\text{Rev}({\pi}(I))$ as the expected revenue elicited by an online algorithm $\pi$ under the specific arrival instance $I$.
Definition 2.2
The competitive ratio of an algorithm $\pi$ is defined as $\text{CR}(\pi)=\inf_{I\in\mathcal{I}}\frac{\text{Rev}({\pi}(I))}{\text{OPT}(I)}$.
Note that competitive ratio is not equivalent to regret. If the competitive ratio of an online algorithm $\pi$ equals 00, this implies that a sequence exists in which the algorithm $\pi$ fails to generate any revenue. However, generally speaking, it is possible that $\pi$ could perform efficiently on the majority of sequences, and if the arrival process is stochastic, $\pi$ could also incur diminishing regret.
In the next two subsections, we outline the most popular algorithmic structures for RM with stochastic and adversarial arrivals. The pseudo-code for all the algorithms introduced below can be found in the Appendix.
2.2 Popular Algorithm Structures under Stochastic Arrivals
Algorithm DLP-PA. The Deterministic Linear Programming with Probabilistic Assignment algorithm (DLP-PA) solves the deterministic linear programming at the beginning and uses the optimal primal solution to make probabilistic decisions, where the DLP formulation is derived by replacing all random variables with their expected values in the hindsight formulation Eq. (1). In the network RM model, since $\lambda_{i}$ is the likelihood of type $i$ arrival in each time epoch, the total number of type $i$ customers arriving between $[0,T]$ is $\Lambda_{i}(T)\sim\text{Bin}(T,\lambda_{i})$. The DLP is defined as:
$\text{DLP}^{\star}=\max_{\mathbf{x}}\Big{\{}T\sum_{i\in[n]}r_{i}x_{i}\quad\text{ s.t. }\sum_{i\in[n]}\mathbf{A}_{i}x_{i}\leq\frac{\mathbf{m}}{T}\text{, }0\leq x_{i}\leq\lambda_{i},\forall i\in[n]\Big{\}},$ (2)
We denote the optimal solution of Eq. (2) as $\mathbf{x}^{*}$ (assuming it is unique). Algorithm DLP-PA then accepts each type $i$ customer with a probability of $\frac{x_{i}^{*}}{\lambda_{i}}$, given sufficient capacity. The details can be found in Algorithm 5 in Appendix 10. The asymptotic regret of Algorithm DLP-PA for classical network RM is $O(\sqrt{T})$ (Gallego and Van Ryzin 1994, 1997).
Note that for $\alpha<\frac{x^{*}_{i}}{\lambda_{i}}(1-\frac{x^{*}_{i}}{\lambda_{i}})$ Algorithm DLP-PA is unfair to all type $i$ customers. This is because if each type $i$ customer is accepted with a probability of $\frac{x^{*}_{i}}{\lambda_{i}}$, it results in a situation where each type $i$ customer is rejected and the prior customer is accepted with probability $\frac{x^{*}_{i}}{\lambda_{i}}(1-\frac{x^{*}_{i}}{\lambda_{i}})>\alpha$.
Algorithm Resolving DLP (RDLP-PA). The RDLP-PA algorithm addresses DLP’s limitation of not accounting for the demand’s randomness after $t=0$. At a high level, RDLP-PA solves the DLP and employs its optimal primal solution, denoted as $\mathbf{x}^{*}$, to make probabilistic assignments. However, at time point $t^{*}$, the applicability of the initial optimal solution $\mathbf{x}^{*}$ for making decisions may be compromised due to the random noise of the stochastic arrival process and probabilistic decision behavior. Hence, we re-solve the DLP with the remaining capacity and remaining time horizon, and continue making probabilistic assignments by the new optimal solution $\mathbf{\tilde{x}^{*}}$ until the end. For network RM, Reiman and Wang (2008) show that the regret incurred by RDLP-PA is $O(T^{1/3})$. Similarly to the DLP-PA algorithm, the RDLP-PA algorithm is unfair to all type $i$ customers for $\alpha<\min\{\frac{x^{*}_{i}}{\lambda_{i}}(1-\frac{x^{*}_{i}}{\lambda_{i}}),\frac{\tilde{x}^{*}_{i}}{\lambda_{i}}(1-\frac{\tilde{x}^{*}_{i}}{\lambda_{i}})\}$.
Algorithm Static Bid Price Control (S-BPC). In contrast to the DLP-PA and RDLP-PA algorithms, bid price control is a threshold-based policy that makes deterministic decisions for every arriving customer. A Static Bid Price Control (S-BPC) uses a consistent threshold price for each leg, which remains unchanged over time. A request is accepted if the revenue it will provide the supplier exceeds the threshold price and is otherwise rejected.
To obtain the optimal fixed threshold price, Williamson (1992), Talluri and Van Ryzin (1998), Talluri et al. (2004) propose solving the dual problem of the DLP at the beginning, denoting the optimal dual variable as $\mathbf{\theta}^{*}\in\mathbb{R}^{L}$ (assuming uniqueness). When each customer of type $i$ arrives, if the revenue $r_{i}$ exceeds the aggregated bid price $\sum_{j=1}^{L}\theta_{j}^{*}A_{ij}$, the customer is accepted; if not, the customer is rejected. The regret of S-BPC for network RM is $O(\sqrt{T})$.
We note that, once the bid price for a type $i$ is established, every customer of type $i$ is either accepted (if there is sufficient capacity) or rejected. Therefore, if resource capacity exceeds demand, Algorithm S-BPC is inherently fair because it either accepts or rejects all customers of type $i$. However, if the demand exceeds the resource capacity, Algorithm S-BPC’s first-come-first-serve nature leads to unequal treatment between earlier and later customers.
2.3 Popular Algorithm Structures under Adversarial Arrivals
Booking Limit (BL). In the BL algorithm (Williamson 1992), the supplier assigns a fixed quota/booking limit $\mathbf{b}=\{b_{1},b_{2},\ldots,b_{n}\}$ for each customer type. In the asymptotic setting, all booking limits scale with the resource quantity; hence, $b_{i}=\Theta(m)$ for $i\in[n]$. The algorithm accepts any type $i$ customers provided that there is sufficient resource capacity and the number of accepted type $i$ customers has not yet reached $b_{i}$. Beyond this point, all type $i$ customers are rejected. The BL algorithm is not individually fair; for any type $i$, the $b_{i}^{\text{th}}$ type $i$ customer is accepted while the $(b_{i}+1)^{\text{th}}$ type $i$ customer is rejected.
NESTING. The NESTING algorithm (Williamson 1992) is applicable for settings with a single type of resource and where customers can be categorized and ranked based on the revenue $r_{i}$ that they add to the supplier. Unlike the approach of assigning individual booking limits to each customer type (which is suboptimal) the NESTING algorithm assigns booking limits to aggregated categories of customer types. For instance, in a situation with three customer types ranked as $r_{1}>r_{2}>r_{3}$, the algorithm sets a booking limit $b_{1}$ for type 1 customers and $b_{2}$ for the combined group of type 1 and type 2 customers. This method demonstrates superior performance compared to traditional booking limits and is considered optimal in the context of single-resource scenarios. However, a significant limitation of the NESTING algorithm is its inapplicability in multi-resource environments, where ranking customers becomes impossible.
The NESTING framework shares some similarities with the BL algorithm. Initially, $n$ quotas $\mathbf{b}=\{b_{1},b_{2},\ldots,b_{n}\}$ need to be established. Following this, $b_{i}$ is assigned as a quota for the total number of type $i$, $i+1$, $\ldots$, $n$ customers. Much like the BL, the NESTING algorithm also lacks individual fairness owing to its quota-based FCFS structure.
After introducing the prevalent algorithmic frameworks, it is imperative to underscore that, albeit our discourse primarily navigates through the classical network revenue management model for elucidatory convenience, our findings retain their applicability across a broad spectrum of more realistic revenue management models delineated in Section 1.2. This is attributed to our methodological refinement of the aforementioned algorithmic frameworks, thereby rendering these algorithms amenable to addressing a wide range of modern revenue management challenges.
3 The “Grace Period”
In this section, we present the concept of the “grace period” and present the analysis of its incorporation into the First-Come-First-Serve algorithm (FCFS). This will serve as a building block for the GP-Enhanced algorithms that we will follow. FCFS works both in the case of stochastic and adversarial arrival sequences in network revenue management problems and is rather simple: it keeps accepting every arriving customer until the capacity of the requested resources is not enough. The primary sources of unfairness in RM settings (and hence also in the case of FCFS) are: (i) limited resources, and (ii) the inherent structure of the algorithm. Note that when the available resources exceed the total demand, then FCFS is always fair since it accepts all customers; the problem arises when there are limited resources compared to the demand faced.
At a high level, the grace period is a time interval during which the algorithm (i.e., FCFS in this case) decides when to stop accepting customers gracefully using randomness.
Definition 3.1
(Decreasing Grace Period) The decreasing grace period $[t_{1}(i),t_{2}(i)]$ for type $i$ customers is defined as: for any $i(k)$ customer that arrives in between round $t_{1}(i)$ and $t_{2}(i)$, $i(k)$ is accepted with a probability of $1-\alpha$ if customer $i(k-1)$ was accepted, and $i(k)$ is rejected if $i(k-1)$ was rejected.
Note that the grace period is defined per customer type, i.e., in general, each customer type will have a different grace period. The exact mathematical instantiation of the parameters of the grace period (i.e., the interval $[t_{1}(i),t_{2}(i)]$) as well as the fairness parameters $\alpha,\delta$ are algorithm specific. To illustrate this, we mathematically instantiate the grace period for the FCFS algorithm below.
3.1 Decreasing Grace Period in Network Revenue Management under FCFS
To achieve $(\alpha,\delta)$-fairness in network revenue management under FCFS, we can utilize the decreasing grace period as follows: for any $\alpha,\delta$, let $\gamma=\log_{1-\alpha}\delta$, and assign a decreasing grace period $[t_{1}(i),t_{2}(i)]$, where $t_{1}(i)=\inf\{t:\min_{j\in[L]}m_{j}(t)\leq\bar{a}n\gamma\}$ and $t_{2}(i)=T$ to each type $i$ customer within the range of $[n]$. Here, $m_{j}(t)$ denotes the remaining capacity of resource $j$ at time $t$. In other words, in order to address the unfairness arising from resource scarcity, it is sufficient for FCFS to have a common decreasing grace period for all customer types. As a result, for FCFS we will drop the dependence on the customer type and denote the grace period as $[t_{1},t_{2}]$. The following theorem states that, by implementing the decreasing grace period as above, we satisfy $(\alpha,\delta)$-fairness, and the revenue loss in comparison to the FCFS algorithm is bounded.
Theorem 3.2
Assigning a decreasing grace period $[t_{1},t_{2}]$, where $t_{1}=\inf\{t:\min_{j\in[L]}m_{j}(t)\leq\bar{a}n\gamma\}$ and $t_{2}=T$ to each type $i$ customer, where $\gamma=\log_{1-\alpha}\delta$, FCFS is $(\alpha,\delta)$-fair for network revenue management. Moreover, compared to the FCFS algorithm without grace period, the revenue loss is bounded by $\frac{\bar{a}}{\underline{a}}n\gamma\bar{r}$, where $\bar{r}=\sup_{i\in[n]}r_{i}$, $\bar{a}=\sup_{i,j}A_{ij}$, and $\underline{a}=\inf_{i,j}A_{ij}$.
To prove Theorem 3.2, we first introduce a lemma below (see Appendix 11 for proof).
Lemma 3.3
Given an online algorithm $\mathcal{A}$, for any customer type $i\in[n]$, and for any $u$, if $\mathbb{P}\left(i(u+1)\succ_{\mathcal{A}}i(u)\right)\leq\alpha$ and $\mathbb{P}\left(i(u-1)\succ_{\mathcal{A}}i(u)\right)\leq\alpha$, then $\mathcal{A}$ satisfies individual fairness (Definition 1.1); namely, for any $u,v\in[n_{i}]$: $\mathbb{P}\left(i(v)\succ_{\mathcal{A}}i(u)\right)\leq\alpha|u-v|$, with probability at least $1-\delta$. Here, the probability mentioned encompasses the overall environmental constraints, including capacity limits, rather than just the algorithm’s decision to accept or reject.
At a high level, the lemma states that for any algorithm $\mathcal{A}$, if for a customer type $i$ the probability that the $u$-th customer gets rejected while the $(u+1)$-th is accepted is upper bounded by $\alpha$ and the same thing holds for customers $u$ and $u-1$, then the algorithm is $(\alpha,\delta)$-fair.
We highlight the intuition of why the revised FCFS is $(\alpha,\delta)$-fair. In the revised algorithm, every customer arriving during $[1,t_{1}-1]$ is accepted, but starting from the period $t_{1}$, the acceptance probability for each customer type progressively decreases. This decrease is conditioned on the preceding customer’s realization. This implies that if the resources are not depleted, the probability that customers of the same type that arrive one after the other get different treatment is bounded by $\alpha$. At round $t_{1}$, the remaining capacity of each resource $j\in[L]$ is at least $\bar{a}n\gamma$. Given that the stopping time (i.e., the time that the algorithm stops accepting them) for each customer type is a geometric random variable, we can apply standard concentration bounds to demonstrate that, with a high probability, all geometric random variables will stop before $\bar{a}n\gamma$ resource $j$ are exhausted for each $j\in[L]$. This concludes the proof that the revised algorithm is $(\alpha,\delta)$-fair.
To obtain a better understanding about the various parameters describing the revenue loss and the efficiency of our proposed algorithm, note that if $\delta=1/m$, and given that $\underline{a}$, $\bar{a}$, $n$ and $\bar{r}$ are constants, Theorem 3.2 implies that the loss in revenue relative to the FCFS algorithm is bounded by $O(\log m)$. Next, we provide the full proof of Theorem 3.2.
Proof 3.4
Proof of Theorem 3.2 First, we show that by assigning a decreasing grace period $[t_{1}(i),t_{2}(i)]$, where $t_{1}(i)=\inf\{t:\min_{j\in[L]}m_{j}(t)\leq\bar{a}n\gamma\}$ and $t_{2}(i)=T$ to each type $i$ customer, Algorithm FCFS is $(\alpha,\delta)$-fair. By Lemma 3.3, we only need to show that for any customer type $i$, with probability at least $1-\delta$, for any $u\in[n_{i}]$, $\mathbb{P}\left(i(u+1)\succ_{\mathcal{A}}i(u)\right)\leq\alpha$ and $\mathbb{P}\left(i(u-1)\succ_{\mathcal{A}}i(u)\right)\leq\alpha$.
Define event $E$ as $E:=\{\text{no resource is depleted in the time interval }[0,T]\}$. We show that $E$ happens with probability at least $1-\delta$ by showing that the complement of $E$ happens with probability at most $\delta$.
[…]
$\leq\mathbb{P}(\exists i\in[n],\text{ s.t. the number of type }i\text{ customers accepted in the grace period }\geq\gamma)$ (3)
$=1-\mathbb{P}(\exists i\in[n],\text{ s.t. the number of type }i\text{ customers accepted in the grace period }<\gamma)$
$=(1-\alpha)^{\gamma}$ (4)
$=(1-\alpha)^{\log_{1-\alpha}\delta}=\delta,$
where (3) is due to the pigeonhole principle, and (4) is because for each type $i$, the number of type $i$ customers accepted after the grace period starts is a geometric random variable with success probability $\alpha$. Therefore, the cdf is $1-(1-\alpha)^{\gamma}$.
Conditional on event $E$ happening, for any customer $i(u)$ arriving within $[1,t_{1}]$, the algorithm accepts $i(u)$ with probability $1$. For any customer $i(u)$ arriving within $[t_{1},t_{2}]$, by the definition of the grace period, we obtain that $\mathbb{P}\left(i(u+1)\succ_{\mathcal{A}}i(u)\right)\leq\alpha$ and $\mathbb{P}\left(i(u-1)\succ_{\mathcal{A}}i(u)\right)\leq\alpha$.
To complete the proof of the theorem, we need to show that the revenue loss is bounded by $\frac{\bar{a}}{\underline{a}}n\gamma\bar{r}$. In the worst case for the revenue, the algorithm rejects all customers after $t_{1}$. By the definition of $t_{1}$, the remaining units for each $j$ are $\bar{a}n\gamma$. Since $\bar{a}n\gamma$ units of resource can serve at most $\frac{\bar{a}}{\underline{a}}n\gamma$ customers, we have that the revenue loss is at most $\frac{\bar{a}}{\underline{a}}n\gamma\bar{r}$.
3.2 Increasing Grace Period
We conclude this section with a related concept, the increasing grace period.
Definition 3.5
The increasing grace period $[t_{1}(i),t_{2}(i)]$ for type $i$ customers is defined as: for any customer $i(k)$ that arrives within the interval from $t_{1}(i)$ to $t_{2}(i)$, customer $i(k)$ is accepted with probability $\alpha$ if customer $i(k-1)$ was rejected, and customer $i(k)$ is accepted if customer $i(k-1)$ was accepted.
The intuition behind the increasing grace period design is that, after a sequence of customer rejections, when the algorithm is ready to start accepting customers again in order to maintain individual fairness, it should not do so abruptly. Instead, the acceptance probability should incrementally increase, with the degree of increase conditioned on the previous customer’s outcome.
4 Individually Fair RM Algorithms under Stochastic Arrivals
Figure 1: Decisions from Algorithm 1 on type $i$ customers
In this section, we focus on making three algorithm structures mentioned in Section 2.2DLP-PA, RDLP-PA, and S-BPC– individually fair without incurring significant additional regret. Based on the techniques we use for these three algorithms, we also enhance a more advanced algorithm – Bid Price Control with Online Gradient Descent (BPC-OGD) to satisfy the fairness metrics in Appendix 13. Further, as Algorithm DLP-PA is a variant for RDLP-PA (resolve 0 times), we analyze DLP-PA and RDLP-PA together.
4.1 Algorithm DLP-PA and RDLP-PA
Here, we introduce the Grace Period Enhanced RDLP-PA algorithm (GP-Enhanced-RDLP). The regret of the original RDLP-PA algorithm is $O(T^{\beta})$, where $\beta$ is a parameter that depends on the frequency of resolving the DLP. For example, $\beta=1/2$ corresponds to resolving zero times (i.e., Algorithm DLP-PA), while $\beta=1/3$ corresponds to resolving one or more times.
We begin by uniformly partitioning the time horizon $[0,T]$ into $K=T^{1-\beta}$ segments of size $T^{\beta}$. We solve the DLP with the initial capacity vector at $t=0$, and solve the DLP with the remaining capacity vector at $t=t^{*}$, where $t^{*}$ is the re-solving time point of the original RDLP-PA algorithm. In the first time segment ($k=1$), for each type $i$ customer, we sample a random variable $y_{i}\sim\text{Bin}(\lambda_{i}T^{\beta},\frac{x_{i}^{*}}{\lambda_{i}})$. Our objective is to accept approximately $y_{i}$ customers of type $i$ in this interval, akin to the original RDLP-PA algorithm’s acceptance of $\text{Bin}(\lambda_{i}T^{\beta},\frac{x_{i}^{*}}{\lambda_{i}})$ customers of type $i$. However, different from RDLP-PA’s probabilistic allocation, we initially accept $y_{i}-\bar{a}n\gamma-1$ customers of type $i$, where $\gamma=\log_{1-\alpha}\delta$. Subsequently, we implement a decreasing grace period for type $i$ customers until the end of the first interval. According to Theorem 3.2, this ensures the individual fairness metrics while accepting $z_{i}^{(1)}=\Theta(\log T)$ fewer customers of type $i$ compared to Algorithm RDLP-PA.
Starting from the second time segment $(k\geq 2)$, due to the probable rejection of the last few customers in segment $k-1$, an immediate acceptance of customers in segment $k$ is unfair. Thus, we introduce an increasing grace period, progressively shifting from rejection to acceptance. Notably, in segment $k-1$, as $z_{i}^{(k-1)}$ fewer customers of type $i$ are accepted compared to Algorithm RDLP-PA, we compensate by accepting $z_{i}^{(k-1)}$ additional type $i$ customers in segment $k$. This is achieved by drawing $y_{i}\sim z_{i}^{(k-1)}+\text{Bin}(\lambda_{i}T^{\beta},\frac{x_{i}^{*}}{\lambda_{i}})$ and implementing an increasing grace period until the arrival of the $(y_{i}-\bar{a}n\gamma)^{\text{th}}$ customer of type $i$. Subsequently, a decreasing grace period is applied until the conclusion of the time segment.
Lastly, following the methodology introduced in Section 3.1, as a resource approaches depletion, an additional decreasing grace period is introduced to conclude the algorithm. The details can be found in Algorithm 1, and the intuition is illustrated in Figure 1.
Input: Preference matrix A. Arriving rate vector $\mathbf{\lambda}$. Capacity vector m. Resolving time point $t^{*}$. Regret bound of input R-DLP algorithm $O(T^{\beta})$. Fairness parameters $(\alpha,\delta)$. $\gamma=\log_{1-\alpha}\delta$.
Output: Optimal solution vectors $\mathbf{x}^{*}$ and $\tilde{\mathbf{x}}^{*}$.
Initialize: Vector $\mathbf{z}^{(0)}=0$. $K=T^{1-\beta}$.
Solve DLP based on the initial capacity, and denote the optimal primal solution as $\mathbf{x}^{*}$.
for $k\in\{1,2,...,\lfloor t^{*}T^{-\beta}\rfloor\}$ do
       Run Algorithm 2 with $t_{1}=(k-1)T^{\beta}$, $t_{2}=kT^{\beta}$, $\mathbf{z}=\mathbf{z}^{(k-1)}$, $\mathbf{x}^{*}$. Get the output $\mathbf{z}^{(k)}\leftarrow\mathbf{w}$.
Solve DLP based on the remaining capacity and time, and denote the optimal solution as $\tilde{\mathbf{x}}^{*}$.
for $k\in\{\lceil t^{*}T^{-\beta}\rceil,\lceil t^{*}T^{-\beta}\rceil+1,\ldots,K\}$ do
       Run Algorithm 2 with $t_{1}=(k-1)T^{\beta}$, $t_{2}=kT^{\beta}$, $\mathbf{z}=\mathbf{z}^{(k-1)}$, $\tilde{\mathbf{x}}^{*}$. Get the output $\mathbf{z}^{(k)}\leftarrow\mathbf{w}$.
Algorithm 1 GP-Enhanced-RDLP
Input: Arriving rate vector $\mathbf{\lambda}$. Capacity vector m. Time horizon $(t_{1},t_{2})$. Fair parameters $(\alpha,\delta)$. Vector $\mathbf{z}$. Optimal primal solution $\mathbf{x}^{*}$.
Output: Vector $\mathbf{w}$, indicating the sum of rejected type $i$ customers.
Sample $y_{i}\sim z_{i}+\text{Bin}((t_{2}-t_{1})\lambda_{i},\frac{x^{*}_{i}}{\lambda_{i}})$.
if $t_{1}\neq 1$ then
       Give an increasing grace period $[t_{1},o_{i}-1]$ to type $i$ customer, where $o_{i}$ is the time when the $(y_{i}-\bar{a}n\gamma)^{\text{th}}$ type $i$ customer arrives.
Give a decreasing grace period $[o_{i},t_{2}]$ to type $i$ customer.
Return vector $\mathbf{w}$, whose $i^{\text{th}}$ element is the sum of number of rejected type $i$ customers in the increasing grace period $[t_{1},o_{i}-1]$ and in the decreasing grace period $[o_{i},\tilde{o_{i}}]$, where $\tilde{o_{i}}$ is the time when the $y_{i}^{\text{th}}$ type $i$ customer arrives.
for $t\in\{t_{1},t_{1}+1,\ldots,t_{2}\}$ do
       if $\min_{j\in[L]}m_{j}(t)-\bar{a}n\gamma\leq 0$ then
             Give a decreasing grace period $[t,T]$ to all types of customers.
Algorithm 2 Auxiliary Algorithm for Algorithm 1
Theorem 4.1
Given any $\alpha,\delta\in(0,1)$, and any input RDLP-PA algorithm that attains a regret of $O(T^{\beta})$ which resolves $K$ times of LP, Algorithm GP-Enhanced-RDLP: (i) is $(\alpha,\delta)$-fair, and (ii) incurs regret $O(\min\{T^{\beta},K\log 1/\delta\})$. Specifically, for $\delta\geq 1/\left(\text{exp}(T^{\beta}-e^{K}\right)$, the regret is $O(T^{\beta})$. [^4] For the state-of-the-art Algorithm (Bumpensanti and Wang 2020), $K=O(\log\log T)$, and the lower bound of $\delta$ is very close to $\exp(-T^{\beta})$. Recall that $1-\delta$ is the confidence level of the individual fairness metrics, here our result shows that no addition asymptotic regret is incurred if the confidence level is no more than $1-\exp(-T^{\beta})$. For large $T$, the confidence level is very close to $1$.
Proof Sketch.
Here, we provide a proof sketch for $K=1$, where Rev(RDLP-PA)re-solves the LP once. Algorithm GP-Enhanced-RDLP is $(\alpha,\delta)$-fair since it introduces a grace period every time that a decision transitions from acceptance to rejection and vice versa. To confirm that the regret of Algorithm GP-Enhanced-RDLP is $O(T^{\beta})$, we have
Regret $=\text{Rev}^{\star}-\text{Rev}(\pi)=\text{Rev}^{\star}-\text{Rev}(\textsf{RDLP-PA})+\text{Rev}(\textsf{RDLP-PA})-\text{Rev}(\pi)$
$=O(T^{\beta})+\text{Rev}(\textsf{RDLP-PA})-\text{Rev}(\pi),$
[…]
Let $\tau$ be a random variable indicating the stopping time at which some resource becomes exhausted (triggering the termination of the algorithm), with $\tau$ belonging to time segment $k(\tau)$. Ignoring the grace period and accepting $y_{i}$ type $i$ customers in each segment, the expected revenue of Algorithm GP-Enhanced-RDLP almost matches that of Algorithm RDLP-PA for $t\in[0,(k(\tau)-1)T^{\beta}]$.
Furthermore, although the application of both increasing and decreasing grace periods in each time segment may lead to the additional loss of up to $\Theta(\log 1/\delta)$ customers of each type, we accept an equivalent number of customers of each type in the following time segment. By doing this, before $t^{*}$, the expected difference in the number of accepted customers of each type between Algorithm RDLP-PA and Algorithm GP-Enhanced-RDLP stems only from the last time segment before $t^{*}$, amounting to $\Theta(\log 1/\delta)$. Therefore, the expected difference in remaining capacity between Algorithm RDLP-PA and Algorithm GP-Enhanced-RDLP is bounded by $\Theta(\log 1\delta)$ at time $t^{*}$. As such, the disparity in the re-solved optimal solutions at $t^{*}$ between Algorithm RDLP-PA and Algorithm GP-Enhanced-RDLP is negligible. Consequently, we eliminate all revenue loss due to the grace period for all $k(\tau)-1$ time segments. For the time segment $k(\tau)$ of size $T^{\beta}$, we may lose a maximum of $O(T^{\beta})$ in revenue compared to the RDLP-PA algorithm. Hence, we have
[…]
The full proof can be found in Appendix 12. ∎
4.2 Algorithm S-BPC
In the algorithm S-BPC, an FCFS approach is employed for each customer type. Drawing intuition from our treatment of FCFS, to guarantee fairness, we only need to offer a decreasing grace period to those types whose bid prices are greater than the aggregated bid prices. From Theorem 3.2, the extra revenue loss of the fair variant is $O(\log 1/\delta)$ for any $\alpha,\delta\in(0,1)$.
Theorem 4.2
Given any $\alpha,\delta\in(0,1)$, and any input S-BPC algorithm with a regret of $O(\sqrt{T})$, by giving a decreasing grace period $[\min_{j\in[L]}m_{j}-\bar{a}n\gamma,T]$ to each type $i$ customer (for $i$ such that $r_{i}>\sum_{j=1}^{L}\theta_{j}^{*}A_{ij}$): (i) the algorithm is $(\alpha,\delta)$-fair, (ii) incurs regret $O(\min\{\sqrt{T},\log 1/\delta\})$. Specifically, for $\delta\geq\text{exp}(-\sqrt{T})$, the regret is $O(\sqrt{T})$.
5 Individually Fair RM Algorithms under Adversarial Arrivals
In this section, we focus on RM problems with adversarial arrivals, where an adversary dictates both the quantity and the sequence of arriving customers of each type. The two prevalent algorithm structures used in this setting are Booking Limit(BL) and NESTING. Next, we introduce how to make these algorithms individually fair.
5.1 Algorithm BL
Algorithm BL sets quotas for each customer $i$, and then makes decisions in an FCFS manner. Consequently, our method of ensuring individual fairness mirrors the modification of FCFS in Section 3. Algorithm GP-Enhanced-BL provides a decreasing grace period as the number of accepted type $i$ customers approaches the quota $b_{i}$, and extends this decreasing grace period to all customer types if the remaining capacity is almost depleted. Details can be found in Algorithm 3.
Input: Preference matrix $\mathbf{A}$. Capacity vector m. Booking limit $\mathbf{b}$. $\gamma=\log_{1-\alpha}\delta$.
Output: Decision on accepting customers over time.
Initialize: Number of accepted customers $s_{i}=0$, $i\in[n]$.
for $t\in\{1,2,...,T\}$ do
       Observe customer of type $i\in[n]$.
       if $\min_{j\in[L]}m_{j}(t)-\bar{a}n\gamma\leq 0$ then
             Give a decreasing period $[t,T]$ to all types of customers.
       else
             if $s_{i}<b_{i}-\gamma$ then
                   Accept the arriving customer, and set $s_{i}\leftarrow s_{i}+1$.
             else
                   Give a decreasing period $[t,T]$ to type $i$ customers.
Algorithm 3 GP-Enhanced-BL
Theorem 5.1
Given any $\alpha,\delta\in(0,1)$,resource capacity $\mathbf{m}$ where $m_{j}=\Theta(m)$ for $j\in[L]$, and any input BL algorithm that attains a competitive ratio of $C$, GP-Enhanced-BL achieves the following: (i) is $(\alpha,\delta)$-fair, (ii) reaches a competitive ratio of $C-O\left(\frac{\log 1/\delta}{m}\right)$.
Theorem 5.1 implies that GP-Enhanced-BL incurs negligible additional loss asymptotically, even under the worst case instance. The detailed proof can be found in Appendix 14. To execute a competitive ratio analysis, we first observe that, given $b_{i}=\Theta(m)$, if the aggregate number of arrivals is $o(m)$, both the original and the revised booking limit algorithms yield a competitive ratio of $1$ since all customers are accepted.
In the event that the total number of arrivals is $\Omega(m)$, the offline optimal revenue is $\Theta(m)$ as the resource capacity scales with $\Theta(m)$. Additionally, any extra loss incurred by GP-Enhanced-BL in comparison to BL is from the decreasing grace period. In the worst scenario, we may lose at most $(\bar{a}+1)n\gamma=\Theta(\log 1/\delta)$ customers. Hence,
$\inf_{I\in\mathcal{I}}\frac{\text{Rev}(\pi(I))}{\text{OPT}(I)}\geq\inf_{I\in\mathcal{I}}\frac{\text{Rev}(\textsf{BL}(I))-(\bar{a}+1)n\gamma\bar{r}}{\text{OPT}(I)}=C-O\left(\frac{\log 1/\delta}{m}\right).$
5.2 Algorithm NESTING
NESTING ranks the customer types based on the revenue they bring to the supplier. Compared to BL, NESTING also sets quotas and accepts customers in a FCFS manner. Differently, each quota $b_{i}$ is the upper bound of number of customers can be accepted with type $i,i+1,\ldots,n$. As such, we still provide a decreasing grace period as the cumulative number of $i$, $i+1$, $\ldots$, $n$ customers approaches the quota $b_{i}$. The rationale behind this modification aligns precisely with that used in BL. Details can be found in Algorithm 4.
Input: Capacity $m$. Nesting quota $\mathbf{b}$.
Output: Decision on accepting customers over time.
Initialize: Number of accepted customers $s_{i}=0$, $i\in[n]$. $\gamma=\log_{1-\alpha}\delta$.
for $t\in\{1,2,...,T\}$ do
       Observe customer of type $i\in[n]$.
       if $m(t)-n\gamma\leq 0$ then
             Give a decreasing period $[t,T]$ to all types of customers.
       else
             if $\sum_{j=i}^{n}s_{j}<b_{i}-n\gamma$ then
                   Accept the arriving customer, and set $s_{i}\leftarrow s_{i}+1$.
             else
                   Give a decreasing period $[t,T]$ to type $i$ customers.
Algorithm 4 GP-Enhanced-NESTING
Theorem 5.2
Given any $\alpha,\delta\in(0,1)$, and any input NESTING algorithm that attains a competitive ratio of $C$, GP-Enhanced-NESTING: (i) is $(\alpha,\delta)$-fair, (ii) reaches a competitive ratio of $C-O\left(\frac{\log 1/\delta}{m}\right)$.
The proof is almost identical to that of Theorem 5.1 and hence we omit it.
6 Numerical Studies
In this section, we conduct a numerical study to evaluate the efficacy of our grace period-enhanced algorithms, using GP-Enhanced-DLP as a representative example, in boosting revenue compared to baseline algorithm DLP-PA. As discussed in Section 4, DLP-PA is characterized as unfair while GP-Enhanced-DLP is considered fair. In this section, we investigate:
  1. 1.
    the expected cumulative revenue for both algorithms without accounting for the negative impacts of unfairness;
  2. 2.
    the expected cumulative revenue when considering the negative impacts of unfairness.
Data Description To build our semi-synthetic dataset, we use a transnational dataset encompassing all transactions recorded from December 1, 2010, to December 9, 2011, for a UK-based, online retail business operating outside traditional store formats. The company specializes in unique gifts suitable for various occasions and primarily serves a wholesale customer base. The dataset is publicly available at https://www.kaggle.com/datasets/tunguz/online-retail. For our analysis, we extracted the following attributes from the dataset:
  • Description (nominal): The name of the product (item) involved in the transaction.
  • Quantity (numeric): The number of units of each product (item) involved in a transaction.
  • UnitPrice (numeric): The price per unit of the product in GBP.
  • CustomerID (nominal): A unique 5-digit identifying integer assigned to each customer.
The dataset comprises over 500,000 transactions, featuring transactions from approximately 5,000 distinct customers, identified by unique CustomerIDs. Customers are classified into 20 distinct categories based on their purchasing patterns by K-means clustering, where $K=20$ and we cluster according to the Description and Quantity of products purchased. As the Description is a nominal variable, we apply one-hot encoding to convert it into a numeric variable. Additionally, the dataset contains over 4,000 unique items listed under the Description column. To clarify, we perform K-modes clustering, where $K=40$, to group all inventories into 40 distinct product types.
Simulation Setup In this section, we focus on comparing the performances of the DLP-PA Algorithm with the GP-Enhanced-DLP Algorithm (GP-Enhanced-RDLP only solves the DLP once). Initially, we construct the preference matrix A, where each element $A_{ij}$ represents the average quantity of product $j$ purchased by customer type $i$, for $i\in[20]$ (customer types) and $j\in[40]$ (product types). Furthermore, we introduce a reward vector r, with each component $r_{i}$ reflecting the average profit generated by customer type $i$, for $i\in[20]$. We estimate the arrival rate $\lambda_{i}$ for each customer type $i$ by dividing the total count of type $i$ customers by the overall number of customers recorded in the dataset. As the original dataset omits inventory details for each product, we deduce the available units for each product type $j$ by calculating the total consumed units and adjusting this figure with a random scaling factor $q_{j}=\text{Unif}(0.5,0.9)$ for each $j\in[40]$. Utilizing these parameters, we proceed to solve the DLP as defined in equation (2), and we can obtain the optimal primal solution $\textbf{x}^{*}$. We assess the model over a time horizon $T$, which varies from 50,000 to 500,000 in increments of 50,000, to examine the dynamics over different operational scales.
We examine two distinct scenarios in our analysis. In the first scenario, individual fairness is disregarded in terms of its impact on revenue. This means that perceived unfairness towards a customer does not influence the revenue generated. Through this scenario, we aim to evaluate the revenue differences between the two algorithms, supporting our claim that the revenue loss associated with GP-Enhanced-DLP is minimal, as asymptotically suggested in Theorem 4.1.
In the second scenario, we consider a situation where individual fairness directly affects revenue. According to our survey detailed in Section 7, 56% of participants reported experiencing perceived individual unfairness in real. Thus, we model a scenario where a customer who is rejected, while a same type customer immediately before or after them is accepted, perceives this as unfair with a 56% likelihood. Reflecting the consequences of perceived unfairness, we integrate a revenue penalty based on our survey findings: 62% of such customers would stop purchasing from the store, and 69% would not recommend the store to others, signifying potential revenue loss. In this model, we reduce the total revenue by an amount $cr_{i}$ for each type $i$ customer who perceives unfairness, where $c$ represents the projected impact of the loss, essentially the expected frequency of visits by a type $i$ customer. For the purposes of our simulations, we set $c$ to values of $\{0.5,1,2\}$, to cover a range of possible loss intensities.
Results In the first scenario, we disregard individual fairness to assess its impact on revenue outcomes. As depicted in Figure 2, the discrepancy in total revenue between the two algorithms is marginal, adhering to the $O(\sqrt{T})$ boundary established in Theorem 4.1.
Figure 2:
Conversely, in the second scenario, revenue loss is incurred whenever a type $i$ customer perceives unfairness, quantified as $cr_{i}$. The results, illustrated in Figure 3, indicate that GP-Enhanced-DLP consistently outperforms DLP-PA under these conditions. Notably, the disparity in total revenue between the algorithms widens as $c$ (the scale of perceived unfairness) and $T$ (the time horizon) increase. This pattern underscores the significance of individual fairness in retail environments characterized by repeated customer transactions. Specifically, if a customer is expected to make three purchases (illustrated here when $c=2$), the revenue generated by GP-Enhanced-DLP can be nearly double that of DLP-PA when T=500,000, highlighting the critical importance of addressing individual fairness to sustain and enhance revenue in settings where repeat business is common. Besides the numerical results, we can even derive some theory in this setting:
Figure 3:
Proposition 6.1
[…]
$\tilde{R}(\pi)-\tilde{R}(\textsf{DLP-PA})=O(T).$
Proof 6.2
[…]
$R(\textsf{DLP-PA})-R(\pi)=O(\sqrt{T}).$
Since Algorithm DLP-PA accepts each type $i$ customer with probability $\frac{x_{i}^{\star}}{\lambda_{i}}$, for any type $i$ customer, the probability that the customer is rejected but one of the preceding or succeeding type $i$ customer is accepted is $2\frac{x_{i}^{\star}}{\lambda_{i}}\left(1-\frac{x_{i}^{\star}}{\lambda_{i}}\right)$. Therefore, in expectation, each type $i$ customer will cause a revenue loss of $2qcr_{i}\frac{x_{i}^{\star}}{\lambda_{i}}\left(1-\frac{x_{i}^{\star}}{\lambda_{i}}\right)$, which is a positive constant. Therefore, we can obtain
$R(\textsf{DLP-PA})-\tilde{R}(\textsf{DLP-PA})=O(T).$
For Algorithm GP-Enhanced-DLP, in each $\sqrt{T}$ time segments, at most two customers can feel unfair, where one is in the decreasing grace period and the other is in the increasing one. Therefore, the expected revenue loss is at most $2\sqrt{T}qc\bar{r}$. Therefore,
$R(\pi)-\tilde{R}(\pi)=O(\sqrt{T}),$
which implies that
$\tilde{R}(\pi)-\tilde{R}(\textsf{DLP-PA})=O(T).$
[…]
7 Survey
In this section, we delineate the primary findings derived from our survey which was conducted between October and November 2023 on Prolific [^5] https://www.prolific.com/. We collected responses from 150 respondents. The primary goal was to evaluate three hypotheses:
  1. H1.
    When buyers observe “unfair” (defined according to our Definition 1.1) treatment in a store prior to shopping there, their purchasing intent diminishes.
  2. H2.
    A customer who has previously made a purchase from a store and subsequently perceives unfair treatment is likely to exhibit adverse future buying behaviors.
  3. H3.
    Individuals perceive disparate treatments occurring in the near term as more unfair compared to those in the distant past.
Comprehensive demographic data is available in Appendix 9.1. The respondents were all based in the United States, fluent in English, and mostly gender-balanced (about a half of them are identifying as male and about half as female). The majority identified as white. To ensure the quality of the responses, we incorporated three attention checks. Only the data from respondents who correctly answered at least two of these questions were considered valid. From the initial 150 participants, 140 responses met this criterion.
The goal of our survey was to assess individuals’ buying behavior when they observed other customers receiving discounts of roughly 10% but they themselves did not. We recognize the significance of the specific discount value in this context; people’s valuation of money varies, and a discount perceived as negligible by some might not induce feelings of unfairness to them. To address this, we conducted a comparative survey, wherein the context incorporated a more substantial discount (specifically, 30%). This supplementary survey, which also comprised of 150 participants, is documented in Appendix 9.3. In all the scenarios we considered, we posit that there are two stores TrendThread and Stitch&Style which the survey respondent is considering. The products sold by the two places are identical in terms of quality and overall characteristics. The only difference is with regards to the pricing and the discounts that each stores offers.
7.1 Buying Behavior and Observed Pre-Purchased Unfairness
In this section, we seek to validate hypothesis H1. We begin by presenting two distinct scenarios designed to evoke feelings of unfairness before making a purchase among consumers.
The first scenario revolves around a consumer’s interaction with social media posts describing others’ purchasing experiences. Imagine you come across a social media post wherein a user claims to have a $10\%$ discount on a hoodie from the store TrendThread. You promptly visit the online shopping website, but no such discount is extended to you from TrendThread. Stitch&Style is offering the identical hoodie at exactly the same price as TrendThread. We aim to capture whether consumers, upon encountering the absence of a promised discount at TrendThread, opt to transfer their patronage to Stitch&Style for their purchase [^6] A concise version is presented here for clarity, with the detailed questions available in Appendix 9.2.
Figure 4: Distribution of preferences in response to exposure to discount-related posts on social media
As illustrated in Figure 4, $47.9\%$ of the respondents perceive the different treatment as a significant factor influencing their purchasing decisions. Within the $41.5\%$ who opted for No Preference, there emerges a question: is their indifference attributed to the perceived unfairness or the magnitude of the discount ($10\%$)? In Figure 5, of those participants who initially selected No Preference, more than $20\%$ would divert their purchase to an alternative store if faced with a more substantial discount disparity. From this observation, we infer that approximately $30\%$ of the participants remain indifferent to differential treatment. In contrast, close to $60\%$ would reconsider their store preference upon perceiving some form of unfairness prior to purchase.
Figure 5: Impact of $30\%$ discount on initially indifferent consumers (i.e., “No preference” from Fig 4)
The second scenario tests whether perceptions of pre-purchase unfairness can also come from customer reviews associated with a store. Consider two stores offering an identical hoodie at the same price and with equivalent ratings. The first store, TrendThread, displays one negative comment addressing discount unfairness among five positive comments [^7] An auxiliary analysis was conducted to discern whether the placement of this negative comment has any consequential impact. Refer to Appendix 9.2 for details.. Conversely, the second store, Stitch&Style, presents three negative comments to discount unfairness. The focus is to analyze how the volume of comments related to unfair treatments influences consumer shopping behavior.
Figure 6: Left, Right Distribution of preferences for store TrendThread and Stitch&Style respectively.
As depicted in Figure 6, upon encountering few comments indicating unfair treatment, $28.6\%$ of respondents express a positive inclination towards making a purchase, whereas $27.9\%$ demonstrate a negative inclination. However, in the presence of a trend of unfairness in the comments, the proportion of positive respondents diminishes to $14.3\%$, while those with a negative predisposition drastically surge to over $50\%$.
The findings from both scenarios underscore the observation that a pronounced perception of pre-purchase unfairness can significantly influence consumer behavior, leading many consumers to avoid shopping at the store they see as unfair. This verifies our hypothesis H1.
7.2 Buying Behavior and Observed Post-purchased Unfairness
Next, we test hypothesis H2. We propose a scenario where a customer, having made a purchase at the full price, subsequently learns of another individual who got a discount for the same item shortly thereafter, potentially triggering feelings of unfairness. To assess the repercussions of such a scenario on future consumer behavior, we initially investigate whether respondents perceive unfairness in this situation. Subsequently, we probe their likelihood of repeating purchases at the same store and their propensity to recommend the store to others.
Figure 7: Left: Perceived unfairness level (0 = Not at all, 1 = Extremely); Middle: Customer repurchase likelihood (0 = Definitely no, 1 = Definitely yes); Right: Customer recommendation likelihood (0 = Definitely no, 1 = Definitely yes)
In Figure 7, the left histogram depicts the perceived unfairness distribution, with 00 indicating no perceived unfairness and $1$ denoting extreme unfairness. Note that in order to obtain this information, we asked respondents to choose how unfairly treated they feel, by moving a needle. We ran the experiment with different initial placements of the needle, to abstract away from biases. For more details, we refer the reader to Appendix 9.
A significant number of respondents perceive high levels of unfairness, evidenced by an average response of approximately $0.68$. This indicates that post-purchase differential treatment significantly influences perceptions of unfairness. The middle histogram illustrates respondents’ willingness to repurchase from the store, where 00 represents a definite refusal and $1$ a definite intention to repurchase. The mean response, around $0.38$, suggests a reduced likelihood of repeat patronage due to observed disparities in post-purchase treatment. The right histogram, portrays the willingness to recommend the store with 00 indicating an unwillingness to recommend and $1$ the opposite; it reveals a predominant reluctance to recommend among participants. This trend towards non-recommendation may adversely affect the store’s potential revenue. These findings underline the importance of maintaining post-purchase individual fairness as a key factor influencing customer retention and advocacy, with direct implications for revenue streams.
7.3 Continuous Time Metrics of Individual Fairness
In this part, our objective is to validate hypothesis H3. This investigation also serves to corroborate the mathematical definition of individual fairness metrics presented in Definition 1.1. To assess whether individuals are more sensitive to recent disparities in treatment as opposed to those in the more distant past, we propose the following scenario: Imagine encountering two identical posts, each stating that a user got a $10\%$ discount from TrendThread. The only distinction between these posts is their post time: one was created 5 minutes ago, while the other dates back to a month ago.
Figure 8: Distribution of opinions regarding which post makes the respondents feel more unfairly treated
As illustrated in Figure 8, the predominant initial choice among respondents is the 5-minute post, accounting for $33.6\%$ of the selections. Incorporating those respondents who initially chose Both and subsequently indicated that the 5-minute post evoked stronger feelings of unfairness, we find that $46.4\%$ of participants perceive recent different treatments as more unfair than ones set further back in time. With this $46.4\%$ surpassing the proportions of other available choices, our continuous metrics hypothesis is empirically validated.
8 Extension: Price-based Revenue Management
Price-based RM represents another significant class of RM problems Gallego and Van Ryzin (1994, 1997), Talluri and Van Ryzin (2004). In this paradigm, a seller markets one or multiple products over a finite horizon with a given initial resource inventory. Unlike the quantity-based RM, which determines whether to accept or reject an incoming customer, the decision in price-based RM involves setting a price for each arriving customer. Typically, we assume the existence of a bijective mapping between the price and the purchasing probability of each type of customer, whereby a higher price correlates with a lower purchasing probability. The objective is to maximize the total revenue over the given finite time horizon. We first define individual fairness for price-based RM, drawing upon a similar intuition to Definition 1.1, designed for quantity-based RM.
Definition 8.1
Assume the algorithm $\mathcal{A}$ assigns a price $p_{u}^{(i)}$ to the $u^{\text{th}}$ customer of type $i$. An Algorithm $\mathcal{A}$ is $(\alpha,\delta)$-fair if for all $i\in[n]$, $u,v\in[n_{i}]$, where $n$ is the number of customer type and $n_{i}$ is the number of type $i$ customers who will arrive, with probability at least $1-\delta$, we have
$\mathbb{P}(p_{u}^{(i)}\neq p_{v}^{(i)})\leq\alpha|u-v|.$
Traditional methods for solving price-based revenue management, for instance, those proposed by Gallego and Van Ryzin (1994, 1997), use the DLP-PA algorithm structure, while Jasin and Kumar (2012), Wang and Wang (2022) employs the RDLP-PA algorithm structure. These papers offer static or dynamic pricing strategies aimed at minimizing regret under a stochastic arrival setting. It is evident that neither static nor dynamic pricing strategies satisfy the individual fairness metrics as introduced in Definition 8.1. In the case of static pricing, as the capacity depletes, we assign a price of +$\infty$+ ∞ to arriving customers. For dynamic programming, its inherent structure is entirely not individually fair.
First, we employ the concept of a decreasing grace period, adjusting it to suit the price-based context. Specifically, a decreasing grace period to type $i$ customer within the interval $[t_{1},t_{2}]$ means that for every $u^{\text{th}}$ type $i$ customer arriving within the interval $[t_{1},t_{2}]$, we assign them a price $p_{u-1}^{(i)}$ with a probability of $1-\alpha$, and a price of +$\infty$+ ∞ with a probability of $\alpha$. This is analogous to the decreasing grace period for quantity-based RM.
Subsequently, we use the modification of the static pricing algorithm as an example to illustrate how to apply a decreasing grace period to guarantee individual fairness. A static pricing algorithm assigns a price $p^{(i)}$ to all type $i$ customers as long as sufficient resources exist, and assigns a price of +$\infty$+ ∞ to all customers once resources are exhausted. The following theorem summarizes the results and the proof can be found in Appendix 15:
Theorem 8.2
For the price-based revenue management problem where the price $p^{(i)}$ is given to type $i$ customers, by assigning a decreasing grace period $[t_{1}(i),t_{2}(i)]$, where $t_{1}(i)=\inf\{t:\min_{j\in[L]}m_{j}(t)\leq\bar{a}n\gamma\}$ and $t_{2}(i)=T$ to each type $i$ customer, where $\gamma=\log_{1-\alpha}\delta$, the $(\alpha,\delta)$-fair metrics is satisfied. Moreover, compared to the original static pricing algorithm, the revenue loss is bounded by $\frac{\bar{a}}{\underline{a}}n\gamma\bar{p}$, where $\bar{p}=\sup_{i\in[n]}p^{(i)}$.